The force between two electrical charges is 0.4 N. What would be the force if the charges are doubled, and the distance between them increased by 100%?


1
Expert's answer
2021-04-29T10:41:26-0400

Write the equation for force between "normal" charges:


"F=k\frac{q_1q_2}{r^2}=0.4\text N."

Write the equation for the "modified" system of charges:


"F_m=k\frac{(2q_1)(2q_2)}{(2r)^2}=k\frac{q_1q_2}{r^2}=0.4\text N."